solve please and show steps 3x-6y=9

Answer by elima(1433)

Choose values for x and then solve equation for y;
3x-6y=9
x=1
3(1)-6y=9
3-6y=9
-6y=9-3
-6y=6
y=-1
-------------
x=-1
3(-1)-6y=9
-3-6y=9
-6y=9+3
-6y=12
y=-2
:)